OpenStudy (anonymous):

Find the lateral area, total area, and volume for the cylinder with the given measurement. r = 2, h = 4 i need the answers in pi

OpenStudy (anonymous):

what are the formulas for the three items you asked for above, based on the information we have worked on?

OpenStudy (anonymous):

Meaning can you tell me the equation for lateral area, total area and volume?

OpenStudy (anonymous):

Lateral Area: L.A. = ph Since finding the perimeter of the base involves the circumference of a circle, the lateral area formula can be written as: L.A. = 2rh (since the circumference of a circle is 2r) Total Area: T.A. = ph + 2B Again, since the base is a circle the perimeter is actually the circumference of the circle. The formula for total area can be written as: T.A. = 2rh + 2r2 Volume:r2h Notice that the volume of a cylinder is simply the area of the base (a circle represented in the formula by r2) multiplied by the height of the cylinder (h).

OpenStudy (anonymous):

Good, so the Lateral area = 2 pi (2) (4) Area = 2 pi (2sqd) + 2 pi (2) (4) vol = pi (2sqd) (4) what are your answers?

OpenStudy (anonymous):

i dont rlly understand the way they wrote it only u quys lol thats the problem

OpenStudy (anonymous):

Ok, so the lateral area = 2 pi x 2 x 4 = how many pi

OpenStudy (anonymous):

what is 2 times 2 times 4 = to?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

correct, so your lateral area = 16pi

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about the others

OpenStudy (anonymous):

so apply the same thing to the area. 2 pi timee (2 x2) + 2pi times (2 x 4)

OpenStudy (anonymous):

and the volume = pi (2x2) times 4

OpenStudy (anonymous):

so my volume is 16 also

OpenStudy (anonymous):

yes, the vol = 16 pi cubed

OpenStudy (anonymous):

what about the surface area what number did you get?

OpenStudy (anonymous):

idk about the total area

OpenStudy (anonymous):

2 pi times (2 x2) + 2pi times (2 x 4) solve this equation

OpenStudy (anonymous):

i qot 6

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok 2 x 2 x 2 = 8 2 x 2 x 4 = 16 8 + 16 = what?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

do you see how i got this? yes!!!!

OpenStudy (anonymous):

lol yes thank you

OpenStudy (anonymous):

Just remember when you are working with area it is squared and when you are working with Volume it is cubed

OpenStudy (anonymous):

ok